OpenStudy (anonymous):

Solve the equation cos^2 A = sin^2 A + sinA when 0º≤A≤360º. There will me multiple answers. A. 270 B.330 C.150 D.30

OpenStudy (anonymous):

cos^2 A=1-sin^2 A you will find a quadratic of sin A solve that

OpenStudy (anonymous):

That is the part that confuses me. I'm unsure as to how i'd set it up to find the quadratic of sin A

OpenStudy (anonymous):

1-sin^2 A=sin^2 A+sinA 2sin^2 A+sin A-1=0

OpenStudy (anonymous):

I appreciate your help, but I'd be lying if I said I had the slightest clue as to what to do next.

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

try to find out the solution of sinA 2sin^2 A + sinA-1=(2sinA-1)(sin A+1) so, sinA=1/2 , -1 can you do the rest yourself?

OpenStudy (anonymous):

equate to 0 by the way^^

OpenStudy (anonymous):

Yeah, I think so. I just plug it in now right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So for my answer I got 56 thats not even an option fml

OpenStudy (anonymous):

sin 30=1/2 sin 270=-1

OpenStudy (anonymous):

Wow. Thank you. I appreciate it.